# Can we accurately read or write quantum data?

### v1

@@ I Introduction
Therefore, the possible limitations of quantum measurements are of particular interest. Wigner [25, 3], followed by Araki and Yanase [1], realized that additive conservation laws restrict what sharp observables can be accurately measured without being disturbed. Wigner showed that the ideal measurement of the spin along an axis is prevented by the conservation of the total angular momentum along an orthogonal axis. From this result, known as the Wigner-Araki-Yanase (WAY) theorem [14], it seems that we have to choose between inaccuracy and disturbing the system even when it is in an eigenstate of the observable. But Wigner also showed that we can avoid this problem if the measuring device is infinitely large.
@@
Despite Wigner’s well known interest in symmetries, he seemed to use the existence of a conservation law as a mere signature of unitarity. But there seems to be a deeper lesson about symmetry encoded in the WAY theorem. A deep relation between symmetries, measurements, and quantum reference frames was clarified by Loveridge, Miyadera, and Busch [15]. However, the constraints on measurements extend beyond symmetries, as we can see from Ozawa’s result that measurements of non-discrete observables are not repeatable [18], and from Tukiainen’s generalization of the WAY theorem beyond conservation laws [23]. For a review of results about limitations of quantum measurements see Busch [2]. But the full scope of the WAY theorem and its generalizations is still being charted.
Ozawa warned about the possible top-down limitations imposed by the WAY theorem in [19, 20]. Lidar replied that it is already taken into account that quantum computing should be implemented on invariant subspaces of the Hilbert space [13]. More investigations of the limits imposed by conservation laws on quantum computing were reported in [12, 11].
More recently it has been shown that ideal projective measurements require infinite resource costs [9]. Without a proper preparation of the pointer state in the “ready” state, the statistics of the results will be biased. Because of the Third Law of Thermodynamics, infinite resource costs are needed for the preparation of the pointer in a suitable state, and also to ensure that the interaction between the observed system and the pointer achieves the correct correlation.
An apparently unrelated factor is the spectrum of the Hamiltonian – it is considered axiomatic to be bounded from below, i.e.  that all energy eigenvalues are larger than some fixed real value [21]. A reason often invoked is that otherwise a system will presumably decay to arbitrarily low energies [8].
Here I show that if the Hamiltonian is bounded from below, no quantum measurement and no state preparation can be accurate. The main result is proved for pure states in Section §II, and generalized to mixed states in Section §III. Section §IV concludes the article with a brief discussion of the implications of these results.
@@ II The main result
Let the pointer observable of the measuring device be, in terms of projectors $\mathsf{\Pi}_{\xi}$ adding up to the identity,
$\mathsf{Z}=\sum_{\xi}\xi\mathsf{\Pi}_{\xi}.$ (2)
The observables $\mathsf{A}$ and $\mathsf{Z}$ may have continuous spectra, but for simplicity we will use the notations (1) and (2). In any case, we can decompose the continuous spectra into discrete regions and be interested only in discrete differences, since anyway in practice we discretize the results of the measurements anyway. But the proofs from this article can be applied to continuous spectra as well.
@@
$\mathbf{U}_{T}\big{(}\psi(0)\otimes\varphi_{\varnothing}\big{)}=\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T),$ (3)
where each $\varphi_{\lambda}(T)$ is an eigenstate of $\mathsf{Z}$ with eigenvalue $\lambda$. We are interested in two particular kinds of processes, measurements and preparations.
@@
$\mathbf{U}_{T}\big{(}\underbrace{\psi_{\lambda}(0)}_{\text{\makebox[0.0pt]{eigenstate}}}\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T).$ (4)
@@
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\underbrace{\psi_{\lambda}(T)}_{\text{\makebox[0.0pt]{eigenstate}}}\otimes\varphi_{\lambda}(T).$ (5)
@@
There is a time interval $[T,T^{\prime}]$, $T^{\prime}>T$, so that, if the total system $S+M$ remains isolated for $t\in[T,T^{\prime}]$, the pointer states $\varphi_{\lambda}(t)$ remain eigenstates of $\mathsf{Z}$ for the eigenvalue $\lambda$, that is,
@@
for any $\lambda$, where $\Pi_{\lambda}\mathcal{H}_{M}$ is the eigenspace of $\mathsf{Z}$ corresponding to the eigenvalue $\lambda$ of $\mathsf{Z}$, cf.  equation (2).
One may imagine more general scenarios, in which the pointer indicates the state of the observed system at a different time than $t=0$ (as for reading outputs) and $t=T$ (as for reading inputs). All of these will be covered by the proof of the following result:
@@
Let $\mathsf{H}$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\mathcal{H}^{\prime}$ a closed subspace of $\mathcal{H}$, and $\Psi\in\mathcal{H}$. If $\mathsf{H}$ is bounded from below and $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all values of $t$ in an open interval $(T,T^{\prime})$, then $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
@@ III Generalization to mixed states
After the interaction between the measuring device and the observed system, the Projection Postulate (or a decoherence-based approach) leaves us with a mixture of states containing pointer eigenstates of $\mathsf{Z}$ for only one of the eigenvalues $\lambda$. That is, it leaves us with the density operator $\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}\rho(T)\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}^{\dagger}$, where $\mathsf{I}_{S}$ is the identity operator on the Hilbert space of the observed system.
@@
Taking into account that density operators on $\mathcal{H}$ are also vectors in the vector space $\textnormal{Herm}(\mathcal{H}\big{)}$, we apply will Lemma 1 to the following case. As a Hilbert space we use the vector space of positive semi-definite trace-class operators with finite Hilbert-Schmidt norm. Let us denote it by $\widetilde{\mathcal{H}}$. It is a real Hilbert space. Anticipating the potential objection that Lemma 1 was proved for complex, not for real Hilbert spaces, we can take the complexified $\widetilde{\mathcal{H}}\otimes\mathbb{C}$ instead of $\widetilde{\mathcal{H}}$, and extend the Hilbert-Schmidt inner product (16) to a Hermitian scalar product, obtaining a Hilbert space. Then the unitary evolution operators $\mathbf{U}_{t}$, extended by linearity to $\widetilde{\mathcal{H}}\otimes\mathbb{C}$, act unitarily on $\widetilde{\mathcal{H}}\otimes\mathbb{C}$. As the vector $\Psi$ from Lemma 1 we will take $\rho(0)$, which is a vector from the space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$. As a closed subspace $\mathcal{H}^{\prime}$ we take the subspace $\widetilde{\Pi}_{\lambda}\widetilde{\mathcal{H}}\otimes\mathbb{C}$.
@@ IV Discussion
We have seen that if the Hamiltonian is bounded from below, accurate measurements and accurate preparations are impossible. Therefore, accurate reading and writing data, being measurements and preparations, are also impossible. This doesn’t prevent us from getting asymptotically closer to perfect accuracy, just from ever reaching it. This limitation is similar to the other limitations of quantum measurements reviewed in the Introduction, but if the Hamiltonian is bounded from below it is stronger, because it prevents measurement accuracy and non-disturbance independently, and it does not need an additive conserved observable that does not commute with the observable $\mathsf{A}$.
@@
At the same time, we may want to reconsider the assumption that our world’s Hamiltonian is bounded from below, as an alternative way to regain some of the lost certainty. This assumption leads to several results that may be problematic. First, it was shown to forbid the existence of a time operator [21], [24]. Second, it was shown to imply that relativistic wave packets confined to a bounded region of space spread instantaneously in the entire space [10]. Third, it implies that there is always a nonvanishing amplitude that the universe returns to a previous state [22]. Finally, in this article it was shown that if the Hamiltonian is bounded from below, it prevents accurate measurements and preparations. Given that all we know about the world, even its macroscopic properties, are known through observations, which are ultimately quantum, this would make everything we know about the external world to be inaccurate.

### v2

@@ I Introduction
Therefore, the possible limitations of quantum measurements are of particular interest. Wigner [35, 4], followed by Araki and Yanase [1], realized that additive conservation laws restrict what sharp observables can be accurately measured without being disturbed. Wigner showed that the ideal measurement of the spin along an axis is prevented by the conservation of the total angular momentum along an orthogonal axis. From this result, known as the Wigner-Araki-Yanase (WAY) theorem [17], it seems that we have to choose between inaccuracy and disturbing the system even when it is in an eigenstate of the observable.
But, in the same paper, Wigner also showed that we can avoid this problem if the measuring device is infinitely large.
@@
Despite Wigner’s well known interest in symmetries, he seemed to use the existence of a conservation law as a mere signature of unitarity. But there seems to be a deeper lesson about symmetry encoded in the WAY theorem. A deep relation between symmetries, measurements, and quantum reference frames was clarified by Loveridge, Miyadera, and Busch [18]. However, the constraints on measurements extend beyond symmetries, as we can see from Ozawa’s result that measurements of non-discrete observables are not repeatable [25], and from Tukiainen’s generalization of the WAY theorem beyond conservation laws [33]. For a review of results about limitations of quantum measurements, up to that date, see Busch [3]. The WAY theorem was later extended to the case when the conserved quantity has a continuous and unbounded spectrum [15]. The measurability and disturbance parts of the WAY theorem were also extended to unsharp observables (positive operator-valued measures), in the presence of bounded conserved quantities, and necessary conditions for either accurate (in an unsharp, probabilistic sense) or non-disturbing measurements can be achieved [22]. But the full scope of the WAY theorem and its generalizations is still being charted.
Another limitation comes from the fact that the Third Law of Thermodynamic prohibits the preparation of pure states [31, 20]. Recently it has been shown that ideal projective measurements require infinite resource costs [10]. Without a proper preparation of the pointer state in the “ready” state, the statistics of the results will be biased. Because of the Third Law of Thermodynamics, infinite resource costs are needed for the preparation of the pointer in a suitable state, and also to ensure that the interaction between the observed system and the pointer achieves the correct correlation. In [21] ways to avoid this limitation were explored, suggesting that sufficiently unsharp observables may avoid it.
Ozawa warned about the possible top-down limitations imposed by the WAY theorem on quantum computing in [27, 28]. Lidar replied that it is already taken into account that quantum computing should be implemented on invariant subspaces of the Hilbert space [16]. More investigations of the limits imposed by conservation laws on quantum computing were reported in [13, 12].
An apparently unrelated factor is the spectrum of the Hamiltonian, which is axiomatically considered to be bounded from below, i.e.  that all energy eigenvalues are larger than some fixed real value [30]. A reason often invoked is that otherwise a system will presumably decay to arbitrarily low energies [9].
Here I show that if the Hamiltonian is bounded from below, no quantum measurement of a sharp observable, and no state preparation, can be accurate. The main result is proved for pure states in Section §II, and generalized to mixed states in Section §III. Section §IV concludes the article with a brief discussion of the implications of these results.
@@ II The main result
Let the pointer observable of the measuring device be, in terms of projectors $\Pi_{\xi}$ adding up to the identity,
$\mathsf{Z}=\sum_{\xi}\xi\Pi_{\xi}.$ (2)
The observables $\mathsf{A}$ and $\mathsf{Z}$ may have continuous spectra, but for simplicity we will use the notations (1) and (2). In any case, we can decompose the continuous spectra into discrete regions and be interested only in discrete differences, since in practice we discretize the results of the measurements anyway. But the proofs from this article can be applied to continuous spectra as well.
@@
$\mathbf{U}_{T}(\psi(0)\otimes\underbrace{\varphi_{\varnothing}}_{\mathclap{\in\Pi_{\varnothing}\mathcal{H}_{M}}})=\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\underbrace{\varphi_{\lambda}(T)}_{\in\Pi_{\lambda}\mathcal{H}_{M}},$ (3)
where $\varphi_{\varnothing}$ and each $\varphi_{\lambda}(T)$ are eigenstates of the pointer observable $\mathsf{Z}$.
We are interested in two particular kinds of processes, measurements and preparations.
@@
$\mathbf{U}_{T}\big{(}\underbrace{\psi_{\lambda}(0)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T).$ (4)
@@
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\underbrace{\psi_{\lambda}(T)}_{\in\mathsf{P}_{\lambda}\mathcal{H}_{S}}\otimes\varphi_{\lambda}(T).$ (5)
@@
There is a time interval $[T,T^{\prime}]$, $T^{\prime}>T$, so that, if the total system $S+M$ remains isolated for $t\in[T,T^{\prime}]$, the pointer states $\varphi_{\lambda}(t)$ remain eigenstates of $\mathsf{Z}$ for each eigenvalue $\lambda$, that is,
@@
for any $t\in[T,T^{\prime}]$ and any $\lambda$, where $\Pi_{\lambda}\mathcal{H}_{M}$ is the eigenspace of $\mathsf{Z}$ corresponding to the eigenvalue $\lambda$ of $\mathsf{Z}$, cf.  equation (2).
Remark 1 (Generality).
The description of a quantum measurement often involves an interaction Hamiltonian,
$\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}+\mu\mathsf{H}_{S+M},$ (8)
where $\mathsf{H}_{S}$ and $\mathsf{H}_{M}$ are the free Hamiltonians of the observed system and the apparatus, and $\mathsf{I}_{S}$ and $\mathsf{I}_{M}$ are the identity operators on the two Hilbert spaces. The measurement involves turning on the coupling $\mu$ between the observed system and the measuring device, and after the measurement turning it off. This is sometimes assumed to be done by an external system, for example an agent, and makes the total Hamiltonian time-dependent. A description in terms of a time-independent Hamiltonian, as for an isolated system, requires the inclusion of this other system, in addition $S$ and $M$. We can add to our equations a Hilbert space $\mathcal{H}_{\mathcal{E}}$ of the environment. However, we can also consider that the Hilbert space $\mathcal{H}_{M}$ includes the additional system, and in fact the entire laboratory, considered isolated. The strength of the interaction also depends on the distance between the observed system and the apparatus, for example spin measurements involve magnetic fields, and the interaction depends on the distance and the orientation of the observed particle. Here we consider the entire laboratory, as well as the position degrees of freedom of its constituent particles, to be included in the system $M$, and also that the observed system $S$ includes its own position degrees of freedom, so that the entire measurement process is autonomous. The pointer observable is then understood as an observable of the entire system $M$. Therefore, there is no loss of generality in the formulation that we are using here. This formulation covers all kinds of sharp measurements, as long as the “ready” pointer state and the states corresponding to the outcomes are assumed to be pure states. It can be verified that all versions of measurements of sharp observables, as listed for example in [6], are particular cases of the description presented here. In Section §III we will extend the results to mixed states, again in the most general settings for sharp measurements.
One may imagine more general scenarios, in which the pointer indicates the state of the observed system at a different time than $t=0$ (measurements) and $t=T$ (preparations). All of these will be covered by the proof of the following result:
@@
Let $\mathsf{H}$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\mathcal{H}^{\prime}$ a closed subspace of $\mathcal{H}$, and $\Psi\in\mathcal{H}$. If $\mathsf{H}$ is bounded from below and $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi\in\mathcal{H}^{\prime}$ for all values of $t$ in an open interval $(T,T^{\prime})$, then $e^{-\frac{i}{\hbar}t\mathsf{H}}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
@@
Remark 3 (Relation with the Yanase condition).
We didn’t appeal to the Yanase condition, which was proposed to ensure the readability of the measurement’s result [1, 26]. This condition requires the apparatus’ part of the additive conserved observable from the WAY theorem to commute with the pointer observable. The reason is that Assumption 1 is more general. For a Hamiltonian as in equation (8), if the coupling becomes $\mu=0$ after the measurement, the Hamiltonian itself is an additive conserved quantity $\mathsf{H}=\mathsf{H}_{S}\otimes\mathsf{I}_{M}+\mathsf{I}_{S}\otimes\mathsf{H}_{M}$, and the Yanase condition becomes
$[\mathsf{Z},\mathsf{H}_{M}]=\mathsf{0}.$ (12)
Therefore, in the case from equation (8), the Yanase condition (12) follows, for a Hamiltonian bounded from below, from Assumption 1 and Lemma 1.
@@ III Generalization to mixed states
After the interaction between the measuring device and the observed system, the Projection Postulate (or a decoherence-based approach) leaves us with a mixture of states containing pointer eigenstates of $\mathsf{Z}$ for only one of the eigenvalues $\lambda$. That is, it leaves us with the density operator $\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}\rho(T)\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}^{\dagger}$.
@@
Taking into account that density operators on $\mathcal{H}$ are also vectors in the vector space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$, we apply will Lemma 1 to the following case. As a Hilbert space we use the vector space of positive semi-definite trace-class operators with finite Hilbert-Schmidt norm. Let us denote it by $\widetilde{\mathcal{H}}$. It is a real Hilbert space. Anticipating the potential objection that Lemma 1 was proved for complex, not for real Hilbert spaces, we can take the complexified $\widetilde{\mathcal{H}}\otimes\mathbb{C}$ instead of $\widetilde{\mathcal{H}}$, and extend the Hilbert-Schmidt inner product (18) to a Hermitian scalar product, obtaining a Hilbert space. Then the unitary evolution operators $\mathbf{U}_{t}$, extended by linearity to $\widetilde{\mathcal{H}}\otimes\mathbb{C}$, act unitarily on $\widetilde{\mathcal{H}}\otimes\mathbb{C}$. As the vector $\Psi$ from Lemma 1 we will take $\rho(0)$, which is a vector from the space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$. As a closed subspace $\mathcal{H}^{\prime}$ we take the subspace $\widetilde{\Pi}_{\lambda}\widetilde{\mathcal{H}}\otimes\mathbb{C}$.
@@ IV Discussion
We have seen that if the Hamiltonian is bounded from below, accurate measurements and accurate preparations are impossible. Therefore, accurate reading and writing data, being measurements and preparations, are also impossible. This doesn’t prevent us from getting asymptotically closer to perfect accuracy, just from ever reaching it. This limitation is similar to the other limitations of quantum measurements reviewed in the Introduction, but if the Hamiltonian is bounded from below it is stronger, because it prevents measurement accuracy and non-disturbance independently and for all situations, and it does not need an additive conserved observable that does not commute with the observable $\mathsf{A}$.
@@
Given that all that we know about the world, even its macroscopic properties, we know through observations, which are ultimately quantum, this would make everything we know about the external world to be inaccurate. All we do in the world, even macroscopically, ultimately rely on quantum preparations, making our actions imprecise. But is this really so? We certainly have very complex electronic devices, including programmable classical computers, and they seem to work very well. They also seem scalable, at least so far we didn’t encounter any limitation of the memory they can have. While they are classical computers, they are still quantum systems, and not only because transistors employ quantum effects, but because the world is quantum. If the existence of classical computers contradicts our results, could it be because the Hamiltonian is actually unbounded? We don’t have a quantitative prediction of the possible accuracy in terms of the size or other properties of the measurement device, so it would be rushed to draw some conclusions from this. But we definitely need to explore this and find out.
At the same time, we may want to reconsider the assumption that our world’s Hamiltonian is bounded from below, as an alternative way to regain some of the lost certainty. This assumption leads to several results that may be problematic. First, it was shown to forbid the existence of a time operator [30], [34]. Second, it was shown to imply that relativistic wave packets confined to a bounded region of space spread instantaneously in the entire space [11], [19]. Third, it implies that there is always a nonvanishing amplitude that the universe returns to a previous state [32]. Finally, in this article it was shown that if the Hamiltonian is bounded from below, it prevents accurate measurements and preparations.
Only time will tell how to solve this dilemma. To avoid closing on a pessimistic note regarding accurate measurements and preparations, even if it turns out that they limit the possibility of scalable quantum computers, we have to remember what we learned from the history of quantum mechanics. Other results, initially perceived as limitations, are behind the incredible development of quantum technologies and other applications: results like complementarity, Bell’s theorem [2], the Kochen-Specker theorem [14], the no cloning theorem [29, 36] and so on, in the hands of visionary thinkers turned out to provide the revolutionary power of quantum technologies. Maybe these results rejected some of our initial intentions, but at the same time they opened unexpected avenues.
Acknowledgments. The author thanks Mohammad Hamed Mohammady for valuable comments and suggestions offered to a previous version of the manuscript. Nevertheless, the author bears full responsibility for the article.